### v1

@@ Abstract
Temporal modeling on regular respiration-induced motions is crucial to image-guided clinical applications. Existing methods cannot simulate temporal motions unless high-dose imaging scans including starting and ending frames exist simultaneously. However, in the preoperative data acquisition stage, the slight movement of patients may result in dynamic backgrounds between the first and last frames in a respiratory period. This additional deviation can hardly be removed by image registration, thus affecting the temporal modeling. To address that limitation, we pioneeringly simulate the regular motion process via the image-to-video (I2V) synthesis framework, which animates with the first frame to forecast future frames of a given length. Besides, to promote the temporal consistency of animated videos, we devise the Temporal Differential Diffusion Model to generate temporal differential fields, which measure the relative differential representations between adjacent frames. The prompt attention layer is devised for fine-grained differential fields, and the field augmented layer is adopted to better interact these fields with the I2V framework, promoting more accurate temporal variation of synthesized videos. Extensive results on ACDC cardiac and 4D Lung datasets reveal that our approach simulates 4D videos along the intrinsic motion trajectory, rivaling other competitive methods on perceptual similarity and temporal consistency. Codes will be available soon.
@@ 1 Introduction
Existing methods utilize flow-based interpolation models [guo2020spatiotemporal, kim2024data] to deduce the regular motion process with the predefined starting and ending frames. And some researchers aim to synthesize the whole videos by means of diffusion models conditioning on these two prompting frames [kim2022diffusion, ho2022video]. Basically speaking, all these methods could barely simulate breathing-induced motions only if two prompting frames exist simultaneously in the test-time evaluation. However, in the preoperative data acquisition stage, the slight movement or unstable breathing of patients may result in dynamic backgrounds between the first and last frames in a respiratory period. This additional bias cannot be thoroughly removed through image registration, thereby affecting the temporal motion modeling. Moreover, clinical research indicates that two high-dose CT or MRI scans will cause prolonged radiation exposure, potentially impairing patients’ health [power2016computed].
@@ 2.3 Network Details of Mo-Diff
Prompt Attention Layer (PAL): To integrate the prompt information into the first-stage TDDM, we do not simply concatenate the first frame with noisy differential fields for training as what previous works did [xing2024cross, peng2023generating]. Instead, a feature-level fusion is adopted to synthesize temporal differential fields with fine-grained textures and shapes, which hold importance for the synthesis performance of future frames. Thus, we propose the PAL to boost the representation of TDDM. Specifically, given the initial frame $\mathcal{I}_{1}$, we define the convolutional mapping $\mathcal{M}$ which can transform the image space into feature space. As visualized in Fig. 2, $\mathcal{M}$ is a combination of convolution, ReLU, and group normalization operations. Then the warped feature of the prompting frame is utilized to enhance features corresponding to temporal differential fields via the element-wise multiplication and addition.

### v2

@@ Abstract
Temporal modeling on regular respiration-induced motions is crucial to image-guided clinical applications. Existing methods cannot simulate temporal motions unless high-dose imaging scans including starting and ending frames exist simultaneously. However, in the preoperative data acquisition stage, the slight movement of patients may result in dynamic backgrounds between the first and last frames in a respiratory period. This additional deviation can hardly be removed by image registration, thus affecting the temporal modeling. To address that limitation, we pioneeringly simulate the regular motion process via the image-to-video (I2V) synthesis framework, which animates with the first frame to forecast future frames of a given length. Besides, to promote the temporal consistency of animated videos, we devise the Temporal Differential Diffusion Model to generate temporal differential fields, which measure the relative differential representations between adjacent frames. The prompt attention layer is devised for fine-grained differential fields, and the field augmented layer is adopted to better interact these fields with the I2V framework, promoting more accurate temporal variation of synthesized videos. Extensive results on ACDC cardiac and 4D Lung datasets reveal that our approach simulates 4D videos along the intrinsic motion trajectory, rivaling other competitive methods on perceptual similarity and temporal consistency. Codes are available at https://github.com/AlexYouXin/Mo-Diff
@@ 1 Introduction
Existing methods utilize flow-based interpolation models [guo2020spatiotemporal, kim2024data] to deduce the regular motion process with the predefined starting and ending frames. And some researchers aim to synthesize the whole videos by means of diffusion models conditioning on these two prompting frames [kim2022diffusion, ho2022video]. Basically speaking, all these methods could barely simulate breathing-induced motions only if two prompting frames exist simultaneously in the test-time evaluation. However, in the preoperative data acquisition stage, the slight movement or unstable breathing of patients may result in dynamic backgrounds between the first and last frames in a respiratory period. This additional bias cannot be thoroughly removed through image registration, thereby affecting the temporal motion modeling. Moreover, clinical research indicates that two high-dose CT or MRI scans will cause prolonged radiation exposure, potentially impairing patients’ health [power2016computed, you2024slord].
@@ 2.3 Network Details of Mo-Diff
Prompt Attention Layer (PAL): To integrate the prompt information into the first-stage TDDM, we do not simply concatenate the first frame with noisy differential fields for training as what previous works did [xing2024cross, peng2023generating, zhang2024pass]. Instead, a feature-level fusion is adopted to synthesize temporal differential fields with fine-grained textures and shapes, which hold importance for the synthesis performance of future frames. Thus, we propose the PAL to boost the representation of TDDM. Specifically, given the initial frame $\mathcal{I}_{1}$, we define the convolutional mapping $\mathcal{M}$ which can transform the image space into feature space. As visualized in Fig. 2, $\mathcal{M}$ is a combination of convolution, ReLU, and group normalization operations. Then the warped feature of the prompting frame is utilized to enhance features corresponding to temporal differential fields via the element-wise multiplication and addition.